### v1

# MLlm-DR: Towards Explainable Depression Recognition with MultiModal Large Language Models

Abstract.
Automated depression diagnosis aims to analyze multimodal information from interview videos to predict participants’ depression scores. Previous studies often lack clear explanations of how these scores were determined, limiting their adoption in clinical practice. While the advent of LLMs provides a possible pathway for explainable depression diagnosis, current LLMs capable of processing multimodal data lack training on interview data, resulting in poor diagnostic performance when used directly. In this paper, we propose a novel multimodal large language model (MLlm-DR) that can understand multimodal information inputs and supports explainable depression diagnosis. MLlm-DR integrates a smaller LLMs and a lightweight query module (LQ-former). Specifically, the smaller LLMs is designed to generate depression scores and corresponding evaluation rationales. To enhance its logical reasoning for domain-specific tasks while maintaining practicality, we constructed a robust training dataset to fine-tune it. Meanwhile, the LQ-former captures depression-related features from speech and visual data, aiding the model’s ability to process multimodal information, to achieve comprehensive depression diagnosis. Our approach achieves state-of-the-art results on two interview-based benchmark datasets, CMDC and E-DAIC-WOZ, demonstrating its effectiveness and superiority.
Depression Recognition, MultiModal, Large Language Models, Affective Computing, Emotion recognition
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 131publicationmonth: 8ccs: Computing methodologies   Artificial intelligenceccs: Human-centered computing  Human-centered computing
1. Introduction
Interviews are widely regarded as the gold standard for diagnosing depression (gold, ). Clinicians evaluate the severity of symptoms by asking participants questions based on well-established diagnostic criteria, such as the DSM-5, ICD-11, and CCMD-3 (DSM-5, ; ICD-11, ; CCMD-3, ). These symptoms include suicidal tendencies, depressive mood, loss of interest, sleep disorders, among others. By combining the scores for each symptom, clinicians determine the final evaluation of depression. However, this process is often time-consuming and can be affected by the clinicians’ subjective judgment.
In recent years, significant progress has been made in automated depression diagnosis based on interviews (sub-attentional, ; MMFF, ; End-to-end, ). These studies analyze the semantics of dialogues and extract emotional cues from speech and facial expressions, leading to efficient and accurate diagnoses (iifdd, ; multimodal, ; MDDR, ; mago, ). However, a common limitation of these methods is their lack of explainability. While they predict participants’ depression scores using neural networks, they do not provide insights into how these scores are determined. This lack of clear rationale can lead to skepticism among clinicians, which restricts the broader adoption of these methods in clinical practice.
Figure 1. The multimodal large language model (MLlm-DR) is designed for explainable depression recognition. It leverages transcribed text, speech, and visual data from participants’ interview videos to generate depression scores and corresponding evaluation rationales.
Explainable depression diagnosis demands that diagnostic models not only output depression scores but also generate explanations for why such scores are given, thereby enhancing the credibility and acceptability of the diagnostic results. The advent of large language models (LLMs) (qwen2.5, ; llama, ; gpt4, ; ernie, ; glm, ) offers a promising pathway to achieving such explainable diagnostics. Specifically, LLMs excel in multi-turn dialogue tasks (zhaochat, ; Instructerc, ; fado, ; huang, ), enabling them to understand and analyze emotional, contextual, and semantic information within interview dialogue to assess participants’ depression scores. Moreover, their powerful logical reasoning (distilling, ; chain, ; kojima, ) and text generation abilities (T5, ; qwen2.5, ) allow them to generate detailed and coherent rationales based on the depression scores they assign. For example, if the model determines a high depression score, it can provide an explanation that highlights specific statements or content from the dialogue that correspond to depressive symptoms.
While LLMs are effective for understanding dialogue content in explainable depression diagnosis, the diagnostic indicators for depression are not limited to participants’ self-reported conversational content alone. They also include objective information, such as the patients’ voice and visual cues (tmac, ; facial, ). These factors are vital as they provide significant insights into an individual’s emotional and cognitive states, thereby offering a more comprehensive set of indicators for depression. Although some existing multimodal LLMs (2023gemini, ; 2024claude, ; Qwen2.5-VL, ) are capable of directly analyzing video data for zero-shot prediction, raw videos are often not provided in depression datasets due to patient privacy constraints, rendering these models inapplicable in real-world scenarios.
In this paper, we propose a novel multimodal large language model (MLlm-DR) that can understand multimodal inputs and supports explainable depression diagnosis, as shown in Figure  1. MLlm-DR is a strong practical framework that utilizes a smaller LLMs yet maintains exceptional logical reasoning capabilities in domain-specific tasks. To achieve this, we leverage advanced LLMs to generate evaluation rationales based on dialogue content, constructing a robust training dataset. This dataset is then used to fine-tune MLlm-DR, ensuring coherent and reliable evaluation rationales.
To further enhance the capabilities of MLlm-DR, enabling it to process multimodal information beyond text, we introduce a lightweight query module (LQ-former). LQ-former utilizes a set of learnable query vectors to extract depression-related features from speech and visual inputs, mapping these features into a unified text feature space compatible with the LLMs. These processed features are then fed into the frozen LLMs for text generation tasks, thereby training the LQ-former. After the training of LQ-former, we freeze it to preserve the learned feature extraction capability. Subsequently, we fine-tune the LLM using a joint optimization strategy that combines language modeling loss and regression loss, targeting both rationale generation and depression score prediction. This design not only improves the accuracy of depression recognition, but also provides clinically consistent and reliable assessments.
Our contributions can be summarized as follows:
  • We propose a novel MLlm-DR framework for explainable depression diagnosis. To the best of our knowledge, this is the first work to propose a multimodal LLMs specifically designed for depression recognition.
  • We construct an explainable depression assessment training data and fine-tune MLlm-DR, enhancing its logical reasoning capabilities while ensuring practicality.
  • We propose the LQ-former module, enabling LLMs to effectively integrate multimodal information for comprehensive depression diagnosis.
  • We evaluate our method on the CMDC and E-DAIC-WOZ datasets, achieving state-of-the-art results.
2. RELATED WORK
2.1. Multimodal Depression Recognition
Recently, multimodal fusion methods have made valuable progress in many depression recognition tasks. For instance, Wei et al. (sub-attentional, ) design independent attention fusion modules for each PHQ-8 sub-score to extract multi-modal features relevant to specific sub-scores, generating individual sub-scores and ultimately aggregating them for overall depression evaluation. Yuan et al. (MMFF, ) introduce multi-order factor decomposition to extract features from single modalities and their cross-modal combinations, significantly enhancing the representational capacity of multi-modal learning and improving model interpretability through a dynamic weighting mechanism. Jung et al. (HiQuE, ) explicitly model the hierarchical structure of interview questions (primary questions and follow-ups), simulating the diagnostic logic of clinicians while leveraging attention mechanisms to identify critical questions and features, enabling more precise depression detection. However, these methods rely on neural networks to directly predict depression scores without providing the corresponding rationale, which reveals a significant limitation.
2.2. LLMs-based Dialogue Understanding
Recently, large language models (LLMs) have made significant progress in dialogue understanding tasks. For instance, Li et al. (empirical, ) demonstrate the potential of LLMs in Dialogue Relation Extraction (DRE) tasks, showing their superior ability to capture long-span and multi-turn dialogue information, outperforming traditional sequence-based and graph-based methods. Lei et al. (Instructerc, ) investigate the use of LLMs for Emotion Recognition in Conversations (ERC). They pretrain the model on a speaker identification task to capture emotional expression characteristics of different roles in dialogues and fine-tune it through multi-task learning by integrating ERC and emotion influence prediction tasks, enhancing performance. Additionally, Huang et al. (huang, ) propose the Emotion-Cause Reasoning Chain (ECR-Chain) framework, leveraging LLMs to analyze statements in dialogues that trigger target emotions and thereby predict causal emotion entailment (CEE). This framework incorporates cognitive appraisal theory to deeply explore the process of emotion generation, thus providing strong interpretability for reasoning outcomes. These studies collectively demonstrate the tremendous potential of LLMs in advancing dialogue understanding tasks.
2.3. Cross-Modal Semantic Alignment
To enable LLMs to process and understand multimodal inputs, researchers have proposed various cross-modal semantic alignment methods(flamingo, ; cogvlm, ), achieving significant progress in multiple cross-modal language generation tasks, such as visual question answering and image captioning. For example, Li et al. (Q, ) propose a semantic alignment module (Q-Former), which utilizes a set of learnable query vectors to capture key information from visual features. Q-Former is trained through a vision-language matching task to align visual and textual features, enabling LLMs to generate textual descriptions based on visual input. To reduce reliance on paired image-text data, Jian et al. (P, ) training their model on textual data to optimize prompts that guide the generation of language from visual inputs. Visual features are subsequently mapped to these prompts, achieving alignment between vision and language. To enhance the model’s understanding of fine-grained visual information, Lu et al. (MQ, ) capture multi-level image features through three submodules: image tagging, object detection, and semantic segmentation. This enables the LLMs to process more granular visual information. Additionally, approaches (Emotion-LLaMA, ) employ simple linear networks (e.g., MLP) to map non-textual features into the textual embedding space, facilitating modality alignment. These studies provide effective methods to extend text-only LLMs to process multimodal data.
Figure 2. The overall framework of the proposed MLlm-DR method. left shows the LQ-former pre-training part, which aims to extracts depression-related feature representations, comprehensible by LLMs, from visual and audio data. On the right is the LLMs fine-tuning part, where the learned feature representation is concatenated with text instruction embeddings as input to fine-tune the LLMs, which then output depression score and evaluation rationales.
3. Methodology
Our method takes a video clip $C_{i}$ from the interview, corresponding to a specific psychological aspect $i$ (e.g., sleep, appetite) of the participant, as input. Each $C_{i}$ includes audio $A$, visual $V$, and transcribed text $T$ data. The method outputs a score $s_{i}$ reflecting the participant’s depression score in aspect $i$, along with the evaluation rationales. The final depression assessment score $S$ is obtained by summing $s_{i}$ across all psychological aspects, aligning closely with the clinical diagnostic process used by physicians.
As showing Figure  2, the method consists of two stages: LQ-former pre-training and LLMs fine-tuning. In the LQ-former pre-training stage, visual or audio features $F$ are first extracted using pre-extraction modules (Encoder). A learnable query vector $Q$ is then used to learn a fixed-length features representation $O$ from $F$. $O$ is subsequently fed into the LLMs with frozen parameters to perform the score prediction task, thereby training LQ-former. In the LLMs fine-tuning stage, the parameters of the pre-trained LQ-former module are frozen. The audio features $F_{a}$ and visual features $F_{v}$ are then used to extract $O_{a}$ and $O_{v}$, which are concatenated with text instruction embeddings as input to the LLMs. Finally, the LLMs is fine-tuned using a joint optimization strategy that combines language model loss $\mathcal{L}_{\text{lm}}$ and regression prediction loss $\mathcal{L}_{\text{mse}}$, enabling the model to output both depression score and evaluation rationales.
3.1. Data Processing
To enhance the performance of smaller LLMs in depression assessment tasks and enable them to generate more logically consistent evaluation rationales, we utilize advanced LLMs (gpt4, ; llama, ) to generate evaluation rationales based on dialogue content and construct training dataset. Specifically, we provide a text instruction requesting the advanced LLMs to produce an evaluation score within a range of 0-3 along with corresponding rationales based on the dialogue content. In this text instruction, we report the actual evaluation score of the patient, ensuring consistency between the rationales generated by the LLMs and the actual score, as shown in the Table 1. The training dataset is then used to fine-tune MLlm-DR, aiming to transfer the reasoning capabilities of the advanced LLMs to a smaller model for domain-specific tasks. In addition, we use HuBERT (hubert, ) to extract audio features from raw speech data. For visual features, we utilize the deep representations provided in the dataset, which are extracted using OpenFace 2.0 (OpenFace2.0, ) or ResNet 50 (ResNet, ).
Table 1. Prompt Instructions
Role Content
System You are a psychiatrist, assessing the participant’s mental health in certain aspects through a series of questions. A score of 0 means not at all, 1 means several days, 2 means more than half the days, and 3 means nearly every day.
User Given the participant’s self-rating score of {label}, please evaluate the participant’s performance in {aspect} based on the dialogue content. The output format is as follows: Evaluation Result: A numeric value between 0-3. Evaluation Reason: A concise and logical description. Each output must strictly follow this format, avoiding omissions or confusion.
3.2. Lightweight Query Module
The LQ-former is designed to help LLMs process and understand multimodal information. It consists of a Transformer (attention, ) Decoder and a Projection Module based on a fully connected network. The Decoder takes a set of learnable query vectors $Q$ as input, which interacts internally via self-attention mechanisms and with non-textual modality features (such as audio and visual) via cross-attention mechanisms, thereby capturing depression cues within the non-textual modality features, and generating fixed-length feature representations $H$. $H$ are then the Projected to the same dimensionality as the LLM’s textual embeddings through a projection layer, producing the output $O$, as shown in the Equation  1. In our task, we utilize two LQ-former modules to process audio and visual features separately. The resulting features are concatenated with the tokenized textual instruction embeddings and used as inputs to the LLMs. To indicate the positions of the inserted features, we employ two special tokens, <AudioHere>and <VideoHere>, as markers.
(1) $H$ $=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
$O$ $=\text{Projection}(H),$
where $Q\in{\mathbb{R}^{n\times{D}}}$ is the query vector, while $K=V\in{\mathbb{R}^{m\times{D}}}$ represents the sequence of non-textual modality features, $D$ is the embedding dimension, $n$ denotes a fixed sequence length, and $m$ represents the length of the non-textual feature sequence,
3.3. Multi-Task Joint Fine-Tuning
Existing LLMs primarily focus on text generation tasks, and their performance in score prediction is often limited when the provided prompt lacks sufficient example samples. To enhance the model’s performance in scoring prediction tasks, we introduce a joint optimization strategy combining causal language modeling cross-entropy loss ($\mathcal{L}_{\text{lm}}$) and mean squared error loss ($\mathcal{L}_{\text{mse}}$), as shown in the Equation  2.
(2) $\mathcal{L}_{\text{joint}}=\mathcal{L}_{\text{lm}}+\mathcal{L}_{\text{mse}}.$
To process the regression prediction task, we design a Multi-Head Attention Aggregation Network (MAA). MAA takes the final hidden states $X\in\mathbb{R}^{L\times D}$ of the LLMs as input. It first splits $X$ into $h$ subspaces corresponding to $h$ attention heads, resulting in a tensor $x^{\prime}\in\mathbb{R}^{L\times h\times d}$, Each head then computes token-level attention weights $\alpha_{i}\in\mathbb{R}^{L}$ using a shared fully connected layer. These weights are used to aggregate the token representations into a sentence-level embedding. The embedding is passed through a fully connected layer to output the predicted scores $S_{\text{reg}}$, which are used to calculate the regression loss, as shown in the Equation  3. where $L$ is the sequence length, $h$ is the number of attention heads, and $d=D/h$ is the the dimension of each subspace, $W_{i}\in\mathbb{R}^{d\times 1}$ is a trainable parameter for each head.
(3) $\alpha_{i}$ $=\text{softmax}\left(\text{W}_{i}x_{i}^{\prime\top}\right),$
$S_{\text{reg}}$ $=\text{FC}\left(\sum_{i=1}^{h}\left(\alpha_{i}\cdot x_{i}^{\prime}\right)\right).$
By leveraging multi-task learning, we enable mutual reinforcement between tasks, improving the accuracy of regression prediction while generating explainable reasoning for evaluation scores. Additionally, since LLMs may occasionally fail to follow the instructions and produce results in the required format, leading to missing prediction scores, the regression prediction score can serve as a supplementary result, enhancing the model’s generalization ability.
4. Experimental setup
4.1. Datasets
We conduct experiments on two interview-based multimodal depression datasets (CMDC (CMDC, ) and E-DAIC-WOZ (DAIC, )) to evaluate the proposed method. Both datasets provide raw speech and text data, as well as scale-based scoring results. A score $\geq$ 10 indicates that the participant is experiencing severe depression. Additionally, for patient privacy protection, both datasets only provide the extracted visual features. CMDC: A Chinese dataset containing 78 samples (26 depressed patients and 52 healthy individuals), with PHQ-9 used as the corresponding questionnaire. Some subjects did not complete the video recording, resulting in missing visual features. E-DAIC-WOZ: An extended version of the DAIC-WOZ dataset, including 163 training samples, 56 validation samples, and 56 test samples, with PHQ-8 as the corresponding questionnaire. Our model is designed to assess each aspect independently and then aggregate the results to predict the overall depression score. Therefore, we treat each aspect-level response as an individual training instance, resulting in a total of $78\times 9=702$ training samples in CMDC and $163\times 8=1304$ in E-DAIC-WOZ.
In the CMDC, the 12 interview questions have clear correspondences with the 9 aspects of PHQ-9. For example, questions 4 and 6 correspond to the ”loss of interest” aspect, questions 9 and 10 to ”low mood”, and question 8 to ”self-harm or suicidal thoughts”, among others. Therefore, our method selectively uses only the interview content that is clearly associated with each aspect to assess the participant’s score on that aspect. In contrast, in the E-DAIC-WOZ, the interview content primarily consists of a series of open-ended questions and does not have explicit correspondences with the items of PHQ-8. Thus, our method utilizes the complete interview content to assess the participant’s depression score for each aspect.
4.2. Implementation Details
The method use LLaMA-3-8B (llama, ) as the base model with fine-tuning and functional extensions, and GPT-4o (gpt4, ) to construct the training datasets. The speech features are deep representations of 768-dim, while the visual features are deep representations of 709-dim or 2048-dim. For missing sequence features, zero padding is applied. LQ-former uses 32 query vectors, each with a dimension of 768-dim, a hidden layer dimension of 1024-dim, and 4 network layers, with a dropout rate set to 0.3. The projection layer employs a two-layer fully connected network with a hidden layer dimension of 1024-dim and an output dimension of 4096-dim. MAA uses 8 fully connected networks to learn the sequence weights for each attention head, followed by a two-layer fully connected network for regression prediction. We fine-tune the query and value projection matrices ($W_{q}$ and $W_{v}$) using LoRA by setting $r=16$, $\alpha=32$ and dropout = 0.1. The learning rate is set to 0.00001, and the model is trained for 10 epochs. All experiments are conducted using the PyTorch deep learning framework, and the training is performed on two A800 GPUs.
4.3. Metrics
To comprehensively evaluate the model’s performance, we employed a variety of metrics. For binary classification tasks, Precision, Recall, and F1-Score were used to measure the accuracy of classification. For regression tasks, Concordance Correlation Coefficient (CCC), as shown in the Equation  4, Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) were utilized to assess the differences between predicted and actual values. These metrics provide a multidimensional evaluation, offering a comprehensive understanding of the model’s performance.
(4) $\text{CCC}=\frac{2\rho\sigma_{x}\sigma_{y}}{\sigma_{x}^{2}+\sigma_{y}^{2}+(\mu_{x}-\mu_{y})^{2}},$
where $\rho$ represents the Pearson correlation coefficient, $\sigma_{x}$ and $\sigma_{y}$ denote the standard deviations of the ground truth and predicted values, respectively, and $\mu_{x}$ and $\mu_{y}$ are their corresponding means. CCC values range from -1 to 1, with 1 indicating perfect agreement.
5. Experimental Results and Analysis
5.1. Baseline Methods
In our experiments, we introduce the following baseline methods for comparison: 1) CubeMLP(cubemlp, ): Utilizes multilayer perceptrons (MLPs) to perform feature fusion across multiple dimensions, enabling comprehensive interactions between modalities. 2) MulT(mult, ): A cross-attention-based multimodal fusion method designed to model interactions across different modalities. 3) MMFF(MMFF, ): Extracting multi-order factors from different modalities and their combinations, providing richer representational capacity for multimodal learning. 4) IIFDD(iifdd, ): Integrating intra-modality and inter-modality feature fusion frameworks for multimodal depression recognition. 5) HiQuE(HiQuE, ): Enhancing depression diagnosis accuracy by analyzing the importance weights of different questions within each modality. 6) LLaMA-3-8B(llama, ): Directly using the LLaMA-3-8B model to analyze dialogue content and generate evaluation results. 7) LoRA(lora, ): Fine-tuning the LLaMA-3-8B with the LoRA method using the training dataset. 8) P+W+A(Harnessing, ): Using prompts to guide LLMs in extracting textual features from transcripts, which are then fused with other modality features (AU, pose, gaze) for multimodal fusion.
5.2. Comparative Experimental
Table 2. Comparison of the performance of different methods on the CMDC and E-DAIC-WOZ datasets (i.e., {T} indicates that only the text modality was used, {A+T} indicates that both audio and text modalities were used, while others indicate that all modalities were used, including text, audio, and visual data. indicates methods that are based on LLMs).
Data Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC CubeMLP \ 0.38 0.8 0.51 \ \
MulT {A+T} \ 0.72 \ \ 4.59 3.66
MMFF \ 0.83 \ \ 4.29 3.19
IIFDD \ 0.95 0.93 0.93 \ \
LLaMA-3-8B {T} 0.37 0.35 1 0.52 10.19 9.36
LoRA {T} 0.80 1.00 0.83 0.91 4.61 4.08
Ours 0.91 1 1 1 3.10 2.61
E-DAIC- WOZ CubeMLP 0.58 \ \ \ \ 4.37
MulT \ 0.64 0.65 0.64 \ \
MMFF 0.67 \ \ \ 4.91 3.98
HiQuE \ 0.71 0.70 0.70 \ \
LLaMA-3-8B {T} 0.54 0.56 0.85 0.68 5.55 4.23
LoRA {T} 0.69 0.74 0.80 0.77 5.04 4.04
P+W+A \ \ \ \ 4.66 3.86
Ours 0.72 0.77 0.81 0.79 4.59 3.41
We compared our proposed method with the latest in inroaches on the CMDC and E-DAIC-WOZ interview-based depression datasets, as shown in Table  2. The results demonstrate that our method outperforms all existing schemes, achieving state-of-the-art performance. Specifically, on the CMDC dataset, our method achieved an exceptional 100% in Precision, Recall, and F1 Score. On the E-DAIC-WOZ dataset, it also significantly outperformed other methods across all metrics. This superior performance is attributed to the LLMs’s powerful text understanding capabilities, which enable it to analyze the participant’s psychological state in specific aspects based on multi-turn dialogue content. Compared to the P+W+A in inroach (using LLMs), our method retains its advantage by integrating multimodal information and LLMs within a unified framework, reducing potential information loss in staged processing pipelines. Directly using LLaMA-3-8B for prediction does not perform well, primarily due to two factors:1) the limited logical reasoning capabilities of smaller LLMs, which make it challenging to provide accurate evaluations based on dialogue content; 2) the inherent limitations of language models in regression prediction tasks, stemming from a lack of domain-specific knowledge and sufficient labeled sample guidance. However, these issues were significantly mitigated through fine-tuning on training datasets, leading to substantial performance improvements.
5.3. Ablation Study
To verify the effectiveness of the LQ-former module and the multi-task learning strategy, we conduct extensive ablation experiments, as shown in Tables  3 and  4. The experimental settings include: 1) w/o Joint: Uses the LQ-former module to integrate audio and visual information and fine-tunes the LLMs, without multitask learning. 2) w/o LQ: Excludes the LQ-former module, not integrating audio and visual information, but fine-tunes with joint optimization. 3) w/o LQ-A: Excludes the audio LQ-former module but retains the visual LQ-former module, fine-tuning with joint optimization. 4) w/o LQ-V: Excludes the visual LQ-former module but retains the audio LQ-former module, fine-tuning with joint optimization. Experimental results demonstrate that the LQ-former module and the multi-task learning strategy can significantly enhance model performance, and removing either one would result in a performance decline. The performance improvement of w/o Joint is due to the pre-trained LQ-former module, which extracts depression-related features from audio and visual data, enhancing the LLM’s ability to process cross-modal information. The performance improvement of w/o LQ benefits from the mutual promotion between multiple tasks. The introduction of a regression prediction task enables the LLMs to generate scores that more closely align with the true values, even in the absence of sufficient prompt samples. Additionally, removing either audio or visual data in the LQ-former module also leads to a slight decline in model performance.
Table 3. Result of ablation study on the CMDC dataset.
Data Model CCC Pre Rec F1 RMSE MAE
CMDC w/o Joint 0.89 1 1 1 3.58 3.08
w/o LQ 0.87 1 1 1 4.18 3.61
w/o LQ-A 0.87 1 1 1 3.99 3.55
w/o LQ-V 0.90 1 1 1 3.52 2.90
Ours 0.91 1 1 1 3.10 2.61
Table 4. Result of ablation study on the E-DAIC-WOZ dataset.
Data Model CCC Pre Rec F1 RMSE MAE
E-DAIC- WOZ w/o Joint 0.67 0.75 0.71 0.73 4.74 3.58
w/o LQ 0.65 0.67 0.76 0.71 4.86 3.78
w/o LQ-A 0.66 0.70 0.72 0.73 4.80 3.71
w/o LQ-V 0.70 0.75 0.79 0.77 4.55 3.56
Ours 0.72 0.77 0.81 0.79 4.59 3.41
5.4. Analysis of LQ-former
To validate the effectiveness of the features extracted by the LQ-former module from visual and speech data, we report the pre-training results of the LQ-former module, as shown in Tables  5 and  6. The feature representations extracted by LQ-former are fed into a a froze LLMs to generate depression scores, in the format of ”Score: 0”. The quality of the generated results reflects the LQ-former module’s performance in two aspects: first, its ability to extract depression-related features from speech and visual data; second, whether these features can be effectively understood by the LLMs. We did not have the model output the corresponding evaluation rationales, as generating fine-grained explanations solely based on speech and visual information is challenging. The main reason is that speech and visual data lack large-scale, fine-grained emotional labels, which prevents the LLMs from generating detailed and accurate explanations as it would with text data.
We evaluate three configurations: 1) LQ-A: Extract depression-related features from speech data using the LQ-former. 2) LQ-V: Extract depression-related features from visual data using the LQ-former. 3) LQ: Extract depression-related features from both speech and visual data using the LQ-former. The experimental results show that the LQ-former effectively extracts depression-related feature representations from audio and visual data that can be understood by LLMs, with this phenomenon being particularly pronounced in the CMDC dataset. The reason for this difference lies in the task complexity: in the CMDC dataset, the LQ-former extracts depression-related features from specific interview segments (averaging 1 minute), making the task simpler and yielding better performance. In contrast, in the E-DAIC-WOZ dataset, the LQ-former must extract depression-related features from the entire interview content (averaging 20 minutes), which increases the task difficulty and results in poorer performance.
Table 5. Result of LQ-former study on the CMDC dataset.
Data Model CCC Pre Rec F1 RMSE MAE
CMDC LQ-A 0.81 0.94 1 0.97 4.01 3.26
LQ-V 0.29 0.60 0.50 0.55 8.13 6.89
LQ 0.85 1 1 1 3.79 3.06
Table 6. Result of LQ-former study on the E-DAIC-WOZ dataset.
Data Model CCC Pre Rec F1 RMSE MAE
E-DAIC- WOZ LQ-A 0.09 0.45 0.42 0.43 7.81 6.48
LQ-V 0.07 0.50 0.42 0.45 7.88 6.18
LQ 0.22 0.52 0.50 0.51 7.24 6.38
5.5. Analysis of Evaluation Rationales
To further evaluate the interpretability and reliability of the model’s outputs, we conducted a human evaluation involving clinical experts. A total of 100 samples were randomly selected from the test set. Experts were asked to independently score each sample and evaluate the model-generated evaluation rationales corresponding to those samples. Each rationale was rated on a 3-point scale: 3 — Fully agree (the expert would have made the same assessment), 2 — Reasonable (different perspective but similar conclusion), and 1 — Disagree (the rationale was not acceptable). We report four evaluation metrics in Table  7: I) The proportion of model outputs not conforming to the required instruction format; II) Result of expert ratings (RMSE/MAE); III) agreement rate between model predictions and expert ratings; IV) Proportional distribution of expert ratings on the quality of model-generated evaluation rationales (3/2/1). From the results, we draw the following observations: 1) LLaMA-3-8B frequently produced improperly formatted outputs, impairing scoring accuracy, while LoRA fine-tuning effectively mitigated this issue and enhanced model performance. 2) Our method achieved 87% and 73% consistency with expert scores on CMDC and E-DAIC-WOZ respectively, demonstrating strong practical value. 3) Expert ratings also showed limitations, even falling slightly behind our method on E-DAIC-WOZ, reflecting the inherent difficulty of depression assessment and the subtlety of depressive cues that often require additional context to interpret. 4) 77% and 64% of the model’s reasoning was rated as “fully agree” by experts on CMDC and E-DAIC-WOZ respectively, highlighting the interpretability of our approach.
Table 7. Results of Human Expert Evaluations
Methods CMDC E-DAIC-WOZ
LLaMA -3-8B LoRA Ours LLaMA -3-8B LoRA Ours
I 9.26% 0.61% 1.23% 18.52% 4.32% 3.12%
II 0.73 0.85 0.87 0.56 0.64 0.73
III 3.25/2.21 5.01/3.94
IV 0.77/0.14/0.09 0.64/0.19/0.17
5.6. Case Study
To gain deeper insights into the reasoning behaviors of different models, we conducted a qualitative case study on two representative samples from the CMDC and E-DAIC-WOZ datasets, as shown in Figure 3. We focus on three aspects: the consistency between predicted scores and evaluation rationale, the relevance of reasoning to original dialogue, and the logical soundness of the reasoning.
Figure 3. Case analysis of explainable depression recognition on the (a) CMDC and (b) E-DAIC-WOZ datasets. We present the inference results from three different models. “Dialogue content” refers to excerpts from the interview process, “label” represents the participant’s true score, the highlighted orange section represents the model’s predicted score, the highlighted yellow sections indicate the key parts of the dialogue content, and the red text indicates the key explanations in the evaluation rationale that are related to the corresponding scores.
1) LLaMA-3-8B method showed clear deficiencies in score prediction. For instance, although the model recognized that the subject had 8 hours of sleep, it still gave a score of 2 (poor), which did not align with the actual situation. 2) LoRA method fine-tuned model showed improvements, providing a ”mild” reasoning result, but the gap between the predicted scores and actual values remained significant. This limitation is primarily due to the lack of real labels during training, leading to deviations between predicted and actual scores. 3) Our method demonstrated a clear advantage in capturing subtle cues in the dialogue that are difficult to detect, enabling accurate predictions. This advantage is attributed to the integration of multi-modal and multi-task learning modules, which provide the model with latent information, thereby enhancing the accuracy of score predictions.
5.7. Analysis of Data Collection Methods
Our method shows significant performance differences between the CMDC and E-DAIC-WOZ datasets. In the CMDC dataset, the questions in the dialogue content are specifically designed based on the PHQ-9 scale, with each question having a clear correspondence to the participant’s specific psychological state. We leverage this correspondence to select content related to the scale’s questions and assess the participant’s corresponding psychological state. This approach aligns closely with the depression diagnosis process used by clinicians based on interviews and has achieved outstanding results. In contrast, the interview content in the E-DAIC-WOZ dataset is open-ended and lacks a fixed format, significantly increasing the complexity of evaluation. This difference limits the performance of existing models on the E-DAIC-WOZ dataset and highlights the critical role of data collection methods in automated depression recognition. This observation also provides valuable insights into optimizing data collection strategies in this field.
6. Conclusion
In this paper, we propose a novel multimodal large language model (MLlm-DR). The model consists of a smaller LLMs and a lightweight query module (LQ-former), which are designed to generate explainable evaluation rationales and integrate multimodal data, respectively, enabling explainable and comprehensive depression diagnosis. This approach is closely aligned with clinical needs and holds significant practical application value. Our approach achieves state-of-the-art results on two interview-based benchmark datasets (CMDC and E-DAIC-WOZ), demonstrating its effectiveness and superiority. The construction of the training dataset is solely based on text, which may result in the loss of fine-grained emotional information from speech and visual cues, leading to potential bias. In future work, we hope creating a larger-scale, fine-grained depression label set for speech and visual data to further improve our research.